OpenStudy (meilendurcer):

1. Find 3x^2 - y^3 - y^3 - z if x = 3, y = -2, and z = -5. Numerical Answers Expected 2. Find -a^2 - 3b^3 + c^2 + 2b^3 - c^2 if a = 3, b = 2, and c = -3. A. -67 B. -17 C. -31 D. 1

OpenStudy (anonymous):

Substutite the corresponding values into the equation to get the answer. Example Find 5x + 3 if x = 2. 5(2) + 3 = 10 + 3 = 13

OpenStudy (meilendurcer):

@TURITW is number one 10?

OpenStudy (anonymous):

3(3)^2 - (-2)^3 - (-2)^3 - (-5) = 27 + 8 + 8 + 5 = 48

OpenStudy (anonymous):

Remember that - (-2)^3 = 8 because - (-8) = 8 not -8

OpenStudy (meilendurcer):

@TURITW I mean the first question 1. is it 20?

OpenStudy (anonymous):

The 3x^2 - y^3 - y^3 - z question?